# Double Privacy Guard: Robust Traceable Adversarial Watermarking against Face Recognition

### v1

@@ Abstract
The wide deployment of Face Recognition (FR) systems poses risks of privacy leakage. One countermeasure to address this issue is adversarial attacks, which deceive malicious FR searches but simultaneously interfere the normal identity verification of trusted authorizers. In this paper, we propose the first Double Privacy Guard (DPG) scheme based on traceable adversarial watermarking. DPG employs a one-time watermark embedding to deceive unauthorized FR models and allows authorizers to perform identity verification by extracting the watermark. Specifically, we propose an information-guided adversarial attack against FR models. The encoder embeds an identity-specific watermark into the deep feature space of the carrier, guiding recognizable features of the image to deviate from the source identity. We further adopt a collaborative meta-optimization strategy compatible with sub-tasks, which regularizes the joint optimization direction of the encoder and decoder. This strategy enhances the representation of universal carrier features, mitigating multi-objective optimization conflicts in watermarking. Experiments confirm that DPG achieves significant attack success rates and traceability accuracy on state-of-the-art FR models, exhibiting remarkable robustness that outperforms the existing privacy protection methods using adversarial attacks and deep watermarking, or simple combinations of the two. Our work potentially opens up new insights into proactive protection for FR privacy.
I Introduction
Face Recognition (FR) systems are widely used in criminal investigations, security surveillance, and online monitoring, providing significant convenience to social networks. The effectiveness of FR systems also poses a threat to privacy protection [1]. When Personal Users (PUs) upload the facial image to social networks, Trusted Authorizers (TAs) (e.g., network administrators and law enforcement agencies) can use the FR system to authenticate facial image. However, Unauthorized Adversaries (UAs) may also use open source FR systems for malicious FR search of PUs’ face images, leading to user privacy leakage. Therefore, there is an urgent need to develop a comprehensive facial image privacy protection method that safeguards PUs’ facial images from malicious recognition search by UAs without affecting legitimate authentication processes, as illustrated in Figure 1.
One countermeasure to address privacy security concerns is adversarial attacks, which deceive FR models by introducing visible patches [2, 3] or invisible perturbations [4, 5] in images. However, these methods exhibit certain drawbacks, including poor black-box transferability, lack of robustness, and significant semantic modifications to protected images [6], thereby affecting their practical applications in social network scenarios. Most importantly, adversarial attacks irreversibly disrupt the recognizable features of facial images. While this prevents malicious recognition by UAs, it also hinders normal identity verification by TAs, potentially leading to security and regulatory vulnerabilities.
Figure 1: Illustration of DPG protection scenario. Users can upload the encoded and protected facial image to social networks, deceiving unauthorized malicious face recognition. Meanwhile, TAs can perform identity authentication on the image for security investigation.
@@ I Introduction
Adversarial attacks in the privacy protection scenario have prompted an innovative question: Can we transform adversarial examples into verifiable patterns for TAs’ identity verification while deceiving unauthorized FR models? Deep watermarking methods offer a new perspective for verifiable adversarial examples [7], which typically protect the copyright of digital media by invisibly embedding traceable watermarks. The watermark encoder learns the embedding of watermarks in the deep feature space of images to enhance robustness, enabling deployment in complex social network scenarios. However, traditional deep watermark algorithms fail to disrupt malicious face recognition, thus unable to fundamentally address privacy leakage issues.
@@
In this paper, we propose the first Double Privacy Guard (DPG) scheme based on robust and traceable adversarial watermarking. By one-time embedding identity-specific watermarks into source images, DPG deceives unauthorized face recognition while permitting TAs to extract watermarks for identity verification. A comparison of DPG with existing works is shown in Table I. DPG focuses on the following three challenges to achieve comprehensive privacy protection: 1) Traceable and adversarial dual-effectiveness: We introduce an information-guided adversarial attack strategy against FR models. This strategy embeds an identity-specific watermark into the source image using the encoder, guiding the recognizable features of the source image to map onto the target feature domain, deceiving FR models. 2) Transferability to black-box FR models: There is an optimization conflict between the adversariality and invisibility of watermarks. Adversariality enhancements under single-model or basic ensemble overly penalize image quality and lack transferability to black-box FR models. Therefore, we adopt a meta-learning strategy in the multi-task end-to-end optimization, calibrating training errors across distinct tasks, thereby enhancing the encoding of universal carrier features. 3) Model robustness against common image processing: We add a simulated noise pool and design regularization under temporary encoder parameters to adjust the optimization direction of the decoder. This enhances the robustness of the decoder and facilitates the synergistic optimization between the encoder and decoder. Furthermore, multi-task training contributes to the overall robustness of the model.
@@
  • We propose DPG, the first dual-effect privacy protection scheme based on traceable adversarial watermarking, which simultaneously attacks unauthorized FR models while allowing image identity verification by TAs.
@@
  • We conducte extensive experiments on two large facial datasets against six SOTA offline FR models and two online commercial FR systems, validating the significant transferability, traceability and robustness of DPG.
II Related Works
II-A Adversarial Attacks against Face Recognition
Studies have demonstrated the vulnerability of FR models to adversarial examples. White-box attacks [8] and query-based black-box attacks [3] heavily depend on model accessibility, rendering them impractical for real-world applications. Therefore, more privacy protection methods are focusing on black-box attacks. [9] propose AdvHat, a real-world adversarial attack using sticker on the hat. [10] enhance the black-box transferability of adversarial attacks through synthesizing eye makeup patches. To eliminate vulnerabilities arising from low-level pixel modifications and enhance transferability to black-box FR models, some studies design perturbation-based attack methods with PGD [11] or FGSM [12] as underlying optimization algorithms. [13] increase the diversity of surrogate models and obtain ensemble-like effects, significantly improving the transferability. [4] enhances transferability by selecting a attribute recognition task related to FR. However, the perturbation-based methods directly overlay gradients in the pixel domain of the carrier, lacking robustness. In essence, adversarial attacks irreversibly disrupt recognizable features essential for FR, thereby hindering normal identity verification by authorizers and impacting the compliant security oversight.
@@ II-A Adversarial Attacks against Face Recognition
II-B Deep Watermarking
Deep watermarking is extensively employed for copyright tracking in digital media. Recent studies have proposed utilizing deep watermarking methods to enhance the security of facial images. [14] utilize the deep watermarking method to encode watermarks into carrier face images, enabling effective tracing of deeply manipulated facial images. [15] introduce Artificial Fingerprints, which tracks facial images by transferring fingerprints from training data to the deep forgery model. Visible watermarks are employed for annotating copyright information and attacking classification models [16, 17]. However, they are not suitable for privacy protection in facial image scenarios. The watermark encoder deeply embeds watermark information into the robust features of the carrier, and through end-to-end training, it learns the globally optimal solution of the encoder and decoder. This inspires us to design a dual-effect privacy protection scheme for complex social network scenarios on this basis. By defining an information-guided adversarial attack, it can avoid the irrecoverable issue of adversarial perturbations, enabling TAs to decode the watermark for identity verification.
III Threat Model
@@
PUs: Possessing a personal image set $x_{s}\in\mathbb{R}^{i\times h\times w}$ and a unique watermark ${{w_{id}}\in[0,1]}^{1\times l}$ representing personal identity, where $\mathbb{R}$ is a real field of dimension $i\times h\times w$ and $l$ represents the length of the watermark. PUs defaults to unified management by TAs and have no knowledge of the structure and parameters of FR models possessed by UAs.
UAs: They can acquire facial images of PUs on social networks, verifying the identity of the images through open-source FR systems $\mathcal{F}(\cdot)$ to obtain relevant privacy data.
TAs: Represented by security supervision personnel or network administrators, TAs verify the identity of facial images when conducting investigations and collecting evidence.
Users can utilize DPG to protect image privacy. Initially, PUs train the encoder $\mathcal{E}(\cdot)$ and decoder $\mathcal{D}(\cdot)$ of DPG by selecting a target person image $x_{t}\in\mathbb{R}^{1\times h\times w}$ with the different identity for impersonation attact. Before uploading images to social networks, PUs can embed watermarks into images using the trained encoder and authorize the decoder for TAs:
@@
The watermark does not impact the social usability of the carrier, PUs can share the watermarked image $\hat{x}_{s}$ on networks. When UAs perform malicious facial comparisons on $\hat{x}_{s}$, the adversariality of $\hat{x}_{s}$ causes the FR model to confidently identify $\hat{x}_{s}$ and $x_{t}$ as the same identity, thereby protecting the true identity information. Simultaneously, TAs can extract the watermark for identity verification without impacting normal security investigation and evidence collection:
@@
Given the threat scenario, the practical application of DPG necessitates optimization training in three aspects: (1) Invisibility optimization: ensuring the watermark does not significantly affect the visual quality of the carrier; (2) Adversarial optimization: enabling the watermarked image to disrupt unknown FR models; (3) Traceability optimization: ensuring the decoder accurately extracts the watermark. These three optimization objectives can be formalized as follows:
$\left\{\begin{matrix}\underset{\mathcal{E}}{min}(\mathcal{L}(\mathcal{E}(x_{s}),x_{s})),\quad invisibility\ optimization,\\ \underset{\mathcal{E}}{min}(\mathcal{L}(\mathcal{F}(\hat{x}_{s}),\mathcal{F}(x_{t}))),\quad adversarial\ optimization,\\ \underset{\mathcal{D}}{min}(\mathcal{L}(\mathcal{D}(\hat{x}_{s}),w_{id})),\quad traceability\ optimization,\end{matrix}\right.$ (3)
where $\mathcal{L}$ represents the corresponding loss functions, which will be elaborated in the next section.
IV Methodology
In this section, we provide a detailed explanation of the proposed pipeline and the architecture is depicted in Figure 2.
Algorithm 1 DPG Training Framework
@@ IV Methodology
15:     Update $\mathcal{E}(\phi_{en})$ and $\mathcal{D}(\cdot)$;
@@
Figure 2: The whole pipeline of the proposed DPG. The source imgage is first encoded with the original parameter encoder to obtain the watermarked image. The watermarked image then enters the meta-learning module for adversarial training using a temporary encoder and undergoes traceability optimization through the noise-pool-based watermark decoder.
IV-A Information-Guided Adversarial Attack
The DPG encoder is designed to embed identity-specific watermark into the source image, mapping the recognizable features of the source image to the target identity feature domain, to disrupt the decision of FR models. The backbone network of the encoder is designed based on MBRS [18]. We first encode the deep adversarial features of the carrier through the encoder. To seamlessly embed the watermark into the encoded carrier feature map and enhance the fault tolerance of the watermark, we introduce redundancy to the watermark through a two-dimensional convolutional layer with a stride of 2, expanding it to the same dimension as the carrier. As watermark information is connected to the feature map in the channel dimension, we introduce the CBAM module [19] to enhance the CNN-based encoding architecture. By explicitly modeling interdependencies between channels and spatial dimensions, CBAM module adaptively recalibrates feature responses to determine the optimal watermark embedding strength.

### v2

# DIP-Watermark: A Double Identity Protection Method Based on Robust Adversarial Watermark

@@ Abstract
The wide deployment of Face Recognition (FR) systems poses privacy risks. One countermeasure is adversarial attack, deceiving unauthorized malicious FR, but it also disrupts regular identity verification of trusted authorizers, exacerbating the potential threat of identity impersonation. To address this, we propose the first double identity protection scheme based on traceable adversarial watermarking, termed DIP-Watermark. DIP-Watermark employs a one-time watermark embedding to deceive unauthorized FR models and allows authorizers to perform identity verification by extracting the watermark. Specifically, we propose an information-guided adversarial attack against FR models. The encoder embeds an identity-specific watermark into the deep feature space of the carrier, guiding recognizable features of the image to deviate from the source identity. We further adopt a collaborative meta-optimization strategy compatible with sub-tasks, which regularizes the joint optimization direction of the encoder and decoder. This strategy enhances the representation of universal carrier features, mitigating multi-objective optimization conflicts in watermarking. Experiments confirm that DIP-Watermark achieves significant attack success rates and traceability accuracy on state-of-the-art FR models, exhibiting remarkable robustness that outperforms the existing privacy protection methods using adversarial attacks and deep watermarking, or simple combinations of the two. Our work potentially opens up new insights into proactive protection for FR privacy.
Index Terms:
Watermark, Adversarial attack, Face recognition, Face privacy.
1 Introduction
Benefiting from the rapid development of deep learning and cloud storage technologies, Face Recognition (FR) systems have experienced significant development and widely deployed in the fields of criminal investigation, security monitoring, and online supervision, facilitating genuine identity verification and management based on facial images. Unfortunately, the rise of FR technology also poses a threat to user privacy [1, 2, 3, 4, 5]. Unauthorized Adversaries (UAs) can use open-source FR recognition systems to match faces obtained from online platforms, thereby stealing users’ real identities and disclosing private information.
Figure 1: Illustration of DIP-Watermark protection scenario. Before users upload personal images, the platform verifies their real identity and registers an identity-specific watermark. After embedding the watermark, the platform uploads the watermarked image for sharing, while retaining the watermark decoder, which is authorized to regulatory bodies. When malicious FR adversaries attempt FR on the protected user image downloaded from the platform, the adversariality of the image disrupts the FR model, leading to incorrect matching results. In cases where impersonators attempt to use user images for identity theft while using the platform’s services, the platform employs the decoder for identity verification and rejects the request. Meanwhile, TAs can use the authorized decoder for image identity verification when necessary.
Several identity privacy protection schemes based on adversarial attacks have been proposed to address the unauthorized malicious Facial Recognition (FR) issue [2, 6, 7]. These methods generate adversarial facial images by adding imperceptible adversarial perturbations [8, 9] or visible adversarial patches to images [10, 11], which disrupt the decision-making of FR models and lead them to match incorrect facial images, thus protecting real identity information. However, these methods exhibit certain drawbacks, including poor black-box transferability, lack of robustness, and significant semantic modifications to protected images [7]. Most Importantly, these methods also raise additional privacy concerns. Malicious users may exploit adversarial images for identity impersonation, attempting to use platform services under an assumed identity. Since adversarial examples have lost recognizable features, even Trusted Authorizers (TAs) cannot conduct legitimate and compliant identity verification through FR, affecting normal security supervision. Therefore, there is an urgent need for a comprehensive privacy protection technique that, while countering unauthorized malicious FR, ensures compliant identity verification by TAs, thus avoiding the potential threat of identity impersonation.
@@ 1 Introduction
To address the limitations of current identity privacy protection schemes, this paper introduces an innovative problem: Can we transform adversarial examples into verifiable patterns for identity verification of TAs while deceiving FR models of UAs? Deep watermarking methods offer a new perspective for verifiable adversarial examples [12], which typically protect the copyright of digital media by invisibly embedding traceable watermarks. The watermark encoder learns the embedding of watermarks in the deep feature space of images to enhance robustness, enabling deployment in complex social network scenarios. However, traditional deep watermark algorithms fail to disrupt malicious face recognition, thus unable to fundamentally address privacy leakage issues.
In this paper, we propose the first Dual Identity Protection scheme based on robust and traceable adversarial watermarking, termed  “DIP-Watermark.” DIP-Watermark serves as an adversarial deep watermark that protects user identity privacy from the perspectives of disrupting unauthorized malicious FR models and real identity traceability. As illustrated in Fig. 1, when users upload personal facial images to online platforms, the platform embed identity-specific adversarial watermarks in user images. The embedded watermark does not affect the visual appearance or social usability of the image. Malicious adversaries may collect user images embedded with adversarial watermarks without permission and attempt to disclose users’ real identity information through FR systems. However, the adversarial watermark disrupts the FR model so that adversaries get wrong matching information, thus safeguarding user privacy. Meanwhile, in processes such as copyright claims, online supervision, or criminal investigations, authorized platforms can extract the watermark using the watermark decoder to verify the real identity information of the facial image, thereby mitigating the potential threat of identity impersonation. The comparison between DIP-Watermark and existing works is illustrated in Table I.  “Original” denotes the original scenario, indicating that the generated adversarial samples or watermark images have not undergone any image post-processing.  “Transferability” indicates the extent to which the method exhibits transferability.  “Robust” represents the robust scenario, indicating that the generated images have undergone image post-processing. Checkmarks and crosses indicate whether the method has the performance in the corresponding scenarios.
@@
DIP-Watermark focuses on the following three challenges to achieve comprehensive privacy protection: 1) Traceable and adversarial dual-effectiveness: We introduce an information-guided adversarial attack strategy against FR models. This strategy embeds an identity-specific watermark into the source image using the encoder, guiding the recognizable features of the source image to map onto the target feature domain, deceiving FR models. 2) Transferability to black-box FR models: There is an optimization conflict between the adversariality and invisibility of watermarks. Adversariality enhancements under single-model or basic ensemble overly penalize image quality and lack transferability to black-box FR models. Therefore, we adopt a meta-learning strategy in the multi-task end-to-end optimization, calibrating training errors across distinct tasks, thereby enhancing the encoding of universal carrier features. 3) Model robustness against common image processing: We add a simulated noise pool and design regularization under temporary encoder parameters to adjust the optimization direction of the decoder. This enhances the robustness of the decoder and facilitates the synergistic optimization between the encoder and decoder. Furthermore, multi-task training contributes to the overall robustness of the model.
@@
  • We propose DIP-Watermark, the first dual-effect identity protection scheme based on traceable adversarial watermark. It can simultaneously disrupt unauthorized FR models and enable TAs to perform normal image identity verification tasks.
@@
  • We conduct extensive experiments on two large facial datasets against six SOTA offline FR models and two online commercial FR systems, validating the significant transferability, traceability and robustness of DIP-Watermark.
2 Related Works
2.1 Adversarial Attacks against Face Recognition
Studies have demonstrated the vulnerability of FR models to adversarial examples. White-box attacks [13] and query-based black-box attacks [11] heavily depend on model accessibility, rendering them impractical for real-world applications. Therefore, more privacy protection methods are focusing on black-box attacks for FR models. Sharif et al. [14] propose the first patch-based attacks, which restricts the perturbation to eyeglasses and physically realizes attacks on FR models. Zhu et al. [15] synthesize eye makeup patches to attack FR models, and Yin et al. [16] further improve the realism of eye makeup and enhanced black-box transferability. Komkov et al. [6] propose AdvHat, a real-world adversarial attack using sticker on the hat. To eliminate vulnerabilities arising from low-level pixel modifications and enhance transferability to black-box FR models, some studies design perturbation-based attack methods with PGD [17] or FGSM [18] as underlying optimization algorithms. Zhong et al. [19] increase the diversity of surrogate models and obtain ensemble-like effects, significantly improving the transferability. Li et al. [8] enhances transferability by selecting a attribute recognition task related to FR. However, the perturbation-based methods directly overlay gradients in the pixel domain of the carrier, lacking robustness. In essence, within the context of identity privacy protection, adversarial attacks irreversibly disrupt recognizable features essential for FR, thereby hindering normal identity verification by authorizers and impacting the compliant security oversight, potentially exacerbating the threat of identity impersonation.
@@ 2.1 Adversarial Attacks against Face Recognition
2.2 Deep Watermarking
Deep watermarking is extensively employed for copyright tracking in digital media [12, 21, 22]. Zhu et al. [12] proposed the first end-to-end DNN-based watermarking algorithm, which accomplishes watermark embedding and extraction through an autoencoder structure and enhances watermark image quality through a discriminator. Due to complex network processing, image propagation significantly alters the data distribution of watermark information, impacting its recovery. Consequently, to enhance algorithm robustness, some deep watermarking approaches are focusing on simulated noise layer design. Liu et al. [21] proposed a two-stage separable deep watermarking framework, addressing the limitations of traditional end-to-end training for nondifferentiable noise. Jia et al. [23] proposed a mini-batch real and simulated JPEG compression method, randomly changing the noise layer in mini batches to help the model search for optimal solutions in different directions. With increasing concerns about copyright and privacy protection in facial images, recent studies have proposed utilizing deep watermarking methods to enhance the security of facial images. Wang et al. [20] utilize the deep watermarking method to encode watermarks into carrier face images, enabling effective tracing of deeply manipulated facial images. Yu et al. [24] introduce Artificial Fingerprints, which tracks facial images by transferring fingerprints from training data to the deep forgery model. The watermark encoder deeply embeds watermark information into the robust features of the carrier, and through end-to-end training, it learns the globally optimal solution of the encoder and decoder. This inspires us to design a dual-effect privacy protection scheme for complex social network scenarios on this basis. By defining an information-guided adversarial attack, it can avoid the irrecoverable issue of adversarial perturbations, enabling TAs to decode the watermark for identity verification.
3 Threat Model
@@
3.1 Personal Users (PUs)
PUs possess a set of personal images $x_{s}\in\mathbb{R}^{i\times h\times w}$ ($\mathbb{R}$ is a real field of dimension $i\times h\times w$ ) and intend to anonymously share images on public network platforms. PUs are required to register with their real identity before using platform services and have the right to request platforms to handle privacy protection of uploaded images. This is to prevent unauthorized adversaries from matching and retrieving users’ real identity information through FR systems, thereby avoiding privacy breaches. Simultaneously, PUs registering on the platform implicitly authorize trusted entities, including judicial and regulatory bodies, for identity management. This enables compliant identity verification during criminal investigations or image identity copyright disputes.
3.2 Unauthorized Adversaries (UAs)
We define two types of UAs based on potential malicious scenarios encountered during the dissemination of facial images.
  • FR adversaries: These adversaries possess open-source FR models $\mathcal{F}(\cdot)$ and acquire facial images of users from online platforms without authorization. They attempt to retrieve the real identity of the image by face image matching through the FR model and obtain user’s private information.
  • Impersonation adversaries: Malicious users may exploit other users’ protected facial images for identity impersonation, either by registering on platforms with false identities or by uploading others’ images to disseminate misinformation through identity impersonation. This may lead to adverse effects such as financial loss and reputational damage for the original user.
3.3 Trusted Authorities (TAs)
As trusted authorizers for user identity information, public network platforms require authentic information management of facial images. When users need to share facial images anonymously on the platform, privacy processing is required to protect the real identity information of users in the image from unauthorized adversaries. The platform can apply the DIP-Watermark method to watermark the facial images uploaded by users. This requires the platform to pre-train a pair of watermark encoder $\mathcal{E}(\cdot)$ and decoder $\mathcal{D}(\cdot)$ in the background. It is noteworthy that the platform only requires training the DIP-Watermark model once, enabling it to be utilized for embedding and extracting watermarks from images of diverse users with distinct identities. This significantly economizes computational resources of the platform. The platform can use open-source FR models during the training process, but it does not know which FR model the adversary is using; that is, the adversary’s attack process is completely black-box to the platform.
The user is first authenticated before using the platform services, and registers a unique identity-specific watermark ${{w_{id}}\in[0,1]}^{1\times L}$ ($L$ represents the length of the watermark), which is formulated by the platform. The platform then uses the watermark encoder to embed the watermark invisibly into the user image as follows:
@@ 3.3 Trusted Authorities (TAs)
Due to the invisible embedding of the watermark into the deep feature map of the image, this encoding process does not affect the image’s social usability. Simultaneously, this privacy management process does not interfere with the platform’s normal identity verification procedure. When malicious users attempt to deceive the platform’s FR model using adversarial images, and impersonate others to use platform services, the platform can utilize the watermark decoder to extract watermark information $\hat{w}_{id}$ from the image for identity verification, thereby rejecting usage requests from malicious users with mismatched identities. The watermark decoding process is as follows:
@@
Given the threat scenario, the practical application of DIP-Watermark necessitates optimization training in three aspects: (1) Invisibility optimization: Unlike adversarial patches and face anonymization, DIP aims to ensure that watermark embedding does not significantly affect the visual quality of the carrier, maintaining the social usability of the facial image, making it more suitable for practical application requirements; (2) Adversarial optimization: Since in practice, the network platform does not know the FR model used by the FR adversary, the watermarked image must possess adversarial transferability to disrupt unknown FR models; (3) Traceability optimization: While ensuring the above two aspects, DIP also has to ensure that the decoder accurately extracts the watermark. Especially when images undergo various image processing operations during network transmission, maintaining accurate traceability is essential. These three optimization objectives can be formalized as follows:
$\left\{\begin{matrix}\underset{\mathcal{E}}{min}(\mathcal{L}_{en}(\mathcal{E}(x_{s}),x_{s})),\;invisibility\ optimization,\\ \underset{\mathcal{E}}{min}(\mathcal{L}_{adv}(\mathcal{F}(\hat{x}_{s}),\mathcal{F}(x_{t}))),\;adversarial\ optimization,\\ \underset{\mathcal{D}}{min}(\mathcal{L}_{wm}(\mathcal{D}(\hat{x}_{s}),w_{id})),\;traceability\ optimization,\end{matrix}\right.$ (3)
where $\mathcal{L}_{en}$, $\mathcal{L}_{adv}$ and $\mathcal{L}_{wm}$ represent the corresponding loss functions, which will be elaborated in the next section.
4 Methodology
In this section, we provide a detailed explanation of the proposed DIP-Watermark framework. The overall pipeline of DIP-Watermark is illustrated in Fig. 2, with the complete algorithmic flow presented in Algorithm 1.
Algorithm 1 DIP-Watermark Training Framework
@@ 4 Methodology
15:     Update $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$;
@@
Figure 2: The whole pipeline of the proposed DIP-Watermark. The source image is first encoded with the original parameter encoder to obtain the watermarked image. The watermarked image then enters the meta-learning module for adversarial training using a temporary encoder and undergoes traceability optimization through the noise-pool-based watermark decoder.
4.1 Information-Guided Adversarial Attack
Inspired by the deep watermark model [23, 12], we designed a watermark encoder based on a CNN network structure, processing the carrier image and the watermark sequence in parallel. The carrier first passes through a stacked block Conv-ReLU consisting of the convolutional layer and the ReLU activation. The CBAM [25] module is used as a channel-wise and spatial-wise attention mechanism to enhance CNN encoding, adaptively recalibrating feature responses to determine the optimal watermark embedding strength. Simultaneously, the Conv-Relu layer processes the watermark sequence in parallel. To seamlessly embed the watermark into the encoded carrier feature map and enhance the watermark’s robustness, we introduce redundancy to the watermark using a two-dimensional convolutional layer with a stride of 2, expanding it to the same dimension as the carrier. The CBAM module is similarly applied to enhance watermark feature encoding. Finally, the watermark sequence is concatenated with the carrier feature maps, followed by two-dimensional convolutional fusion, resulting in the watermarked image.